40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Jan 31 07:41:23 2011, Aidan, Update, Green Locking, DFD - medm screen Screenshot-3.png
    Reply  Mon Jan 31 12:40:38 2011, rana, joe, Update, Green Locking, DFD - medm screen a.pdf
       Reply  Mon Jan 31 18:25:25 2011, Aidan, Update, Green Locking, DFD - results from the new filters (and running with AWG) DFD-bandwidth_noise_CLP500_to_300Hz.pdfDFD-bandwidth_noise_CLP500_to_7000kHz.pdf
Message ID: 4232     Entry time: Mon Jan 31 12:40:38 2011     In reply to: 4230     Reply to this: 4234
Author: rana, joe 
Type: Update 
Category: Green Locking 
Subject: DFD - medm screen 

This is a plot showing the old filters and the new ones we added this morning.

The new ones have a Cheby for AC coupling below 10 Hz and then a 500 Hz LP after the mixer. The LP frequency has been increased so that we can use this signal in a feedback loop to the ETM with a ~100 Hz UGF.

Attachment 1: a.pdf  20 kB  | Hide | Hide all
a.pdf
ELOG V3.1.3-